OpenStudy (anonymous):

Need help.... How do i convert this information into an index and the forcast the next year Month Year 1 Year 2 Year 3 Year 4 Forecast 1 18,000 45,100 59,800 35,500 2 19,800 46,530 30,740 51,250 3 15,700 22,100 47,800 34,400 4 53,600 41,350 73,890 68,000 5 83,200 46,000 60,200 68,100 6 72,900 41,800 55,200 61,100 7 55,200 39,800 32,180 62,300 8 57,350 64,100 38,600 66,500 9 15,400 47,600 25,020 31,400 10 27,700 43,050 51,300 36,500 11 21,400 39,300 31,790 16,800 12 17,100 10,300 31,100 18,900 Avg.

OpenStudy (anonymous):

May be you have to take average for each month in 3 years and use that as the forecast for the 4th year

OpenStudy (anonymous):

sorry,take average for 4 years

OpenStudy (anonymous):

Vsahal3 how do i do that. this is what i did. yr1 the 3 highest numbers are 83200+72900+57350=213450 then i divide it by 3 213450/3=71150 and this was the average for yr1. I did this for each yr. but from there i am confused. would the index be the average for all four yrs? like this: yr1 71150, yr2 52743, yr3 64630, & yr5 67533

OpenStudy (anonymous):

For example for Month1of year5 it would be (18000+45100+59800+35500)/4=39600 This is done for all the months